Based on a study conducted on residents in the United States, we found their salaries follow the normally distribution with a mean of $48600, and a standard deviation of $14000.
a) If we randomly select 50 residents from the United States, approximately how many of them will have salaries above $70000?
b) If we randomly select 10 residents from the United States, what is the probability that their average salary is between $45000 to $50000?
c) Knowing Jackie’s salary is 17% below the third quartile, what is her actual salary?
1
Expert's answer
2020-12-28T14:46:03-0500

a)

"P(X > 70000) = P(Z > \frac{70000-48600}{14000}) \\

= P(Z> 1.53) \\

= 0.063"

Therefore Expected number of people with salaries above $70000:

"= 0.063 \times 50 = 3.15"

b)

"P(45000< \bar{X} < 50000) = P(\frac{45000-48600}{\frac{14000}{\sqrt{10}}} <Z< \frac{50000-48600}{\frac{14000}{\sqrt{10}}}) \\

= P(-0.81 < Z < 0.32) \\

= P(Z<0.32)-P(Z <-0.81) \\

= 0.6255 -0.2090 \\

= 0.4165"

c) From standard normal tables:

"P(Z < 0.6745) = 0.75"

The third quartile salary:

"= Mean + 0.6745 \times σ \\

= 48600 + 0.6745 \times 14000 \\

= 58043"

Jackie's salary "= 0.83 \times 58043 = 48175.69"